# Where Quantum Complexity Helps Classical Complexity: A Survey

### v2

@@ 4.3. Graph Traversal Problems
One of the most famous problems in graph theory is the traveling salesman problem. However, the extent to which quantum computers can speed up algorithms to resolve this problem is not yet known. Moylett et al. (moylett2017quantum, ) use a quantum backtracking algorithm to a classical algorithm by Xiao and Nagamochi to demonstrate a quadratic quantum speedup when the degree of each vertex is at least 3. Then, to speed up the classical algorithm for when the degree of each vertex is at least 4, authors use similar techniques to speed up the higher degree graphs by reducing the instances (moylett2017quantum, ).
@@
Salehi et al. (salehi2022unconstrained, ) analyze the traveling salesman problem with time windows (TSPTW) in the context of quantum computation. They introduced quadratic unconstrained and higher-order binary optimization formulations as well as the advantages of edge-based and node-based formulations of TSPTW. In addition, the authors conducted experimental realizations on their methods on a quantum annealing device.
@@ 4.4. Maximum Matching on Graphs
The work in (dorn2009quantum, ) introduces quantum algorithms for the following matching problems. The author considers unweighted and weighted graphs with $n$ vertices and $m$ edges. In general graphs, the results in (dorn2009quantum, ) shows that finding a maximal matching takes $\mathcal{O}(\sqrt{nm}\log^{2}n)$. In addition, a maximum matching under the same setting takes $\mathcal{O}(n\sqrt{m}\log^{2}n)$. Considering a different setting, the author proves that finding a maximum weight matching in bipartite graphs is proportional to $\mathcal{O}(n\sqrt{m}N\log^{2}n)$ s.t. $N$ is the largest edge weight. Whilst a minimum weight perfect matching introduces the time complexity of $\mathcal{O}(n\sqrt{nm}\log^{3}n)$.
QA suits problems that can be formulated as QUBO problems. E.g. SAT, graph coloring, and traveling salesman. McLeod and Sasdelli (mcleod2022benchmarking, ) investigate the performance of D-Wave’s 2000Q (2048 qubits) and Advantage (5640 qubits) quantum annealers empirically. In particular, they benchmark the solvers based on solving a specific instance of the maximum cardinality matching problem. They used the results of a previous work that measured the performance of earlier QA hardware from D-Wave. The authors illustrated that Advantage produces optimal solutions to problems of a larger size, compared to 2000Q. Further, the authors investigated the scaling of the diabatic transition probability, by utilizing the Landau-Zener formula. In particular, they considered the structure of the problem and its implications for suitability to QA. Moreover, they introduced a method, exploring the behavior of minimum energy gaps for scalable problems deployed to QAs. The authors showed for their targeted QA problem that the minimum energy gap does not scale as desired. One might ask if this problem is suitable for benchmarking QA hardware. It might lack the necessary nuance to recognize significant performance improvements between generations.
Given an $L\crossproduct L$ array of qubits. Fowler (fowler2013minimum, ) considers the minimum weight perfect matching problem. Provided $(1)$ a modular architecture s.t. there is a finite maximum number of non-functioning qubits in any given connected defective patch $(2)$ gate, leakage, and loss error rates below some set of nonzero values $(3)$ a 2-D array of processors with finite processing power, finite local memory, and a communication line with their nearest neighbor at a finite data rate $(4)$ external memory to store all detection events and matching data. This should last for the duration of a temporally finite quantum algorithm.
@@ 4.5. Maximum-Cut Problem
Boulebnane and Montanaro (boulebnane2021predicting, ) generalized techniques appeared in (farhi2022quantum, ). Applied to different families of sparse random graphs, the authors evaluated the performance of QAOA in the infinite-size limit for the MAX-CUT problem. Their analysis rigorously holds only when $n\rightarrow\infty$ and large degree limit. Still, the analysis provides sufficient initial variational parameters to optimize small-scale, very sparse instances. In addition, the authors introduced a numerical algorithm. It evaluates the performance of QAOA on an average finite-size instance of the SK model. This is where the analytic methods exposed in (farhi2022quantum, ) fail. Nevertheless, the algorithm supports low depth only. Further, their work is unable to address the D-spin model for $D\geq 3$ and $p>1$.
@@ 4.6. Graph Isomorphism
Mondada and Andrés-Martínez (mondada2023subgraph, ) suggested an algorithm for the subgraph isomorphism problem on port graphs. Their algorithm performs pattern matching in quantum circuits for many patterns simultaneously. This is independent of the number of patterns. The running time of the algorithm is linear for the size of the input quantum circuit, after pre-computation, in which the patterns are compiled into a decision tree.
Consider the connected port graphs. Such graphs introduce a property i.e. every edge $e$ incident to $v$ has a label $L_{v}(e)$ unique in $v$. The algorithm suggested by Mondada and Andrés-Martínez (mondada2023subgraph, ) enumerates the subgraph isomorphism problem $H\subseteq G$ for all $m$ pattern matches in time $\mathcal{O}(P)^{P+\frac{3}{2}}\cdot|V(G)|+\mathcal{O}(m)$ s.t. $P$ is the number of vertices of the largest pattern. Also, they can express the bound obtained with the quantum circuit in terms of the maximum number of qubits $N$ and depth $\delta$ of the patterns: $\mathcal{O}(N)^{N+\frac{1}{2}}\cdot\delta|V(G)|\log(\delta)+\mathcal{O}(m)$.
@@ 4.7. Minimum Vertex Cover Problem
Zhang et al. in (zhang2022applying, ) provided a quantum circuit solution scheme. They used an algorithm of quantum approximate optimization. Based on the Ising model quantized by rotation operators and Pauli operators, a quantum Ising equation and Hamiltonians are obtained. Also, a parametric unitary transformation between the original Hamiltonian and the problem Hamiltonian, as a generator, is acquired. The final quantum state and the problem of Hamilton’s expectation are generated by a continuous evolution of two parametric unitary transformations. The results of simulations in (zhang2022applying, ) indicate that the model may solve problems with large probability in linear time. It can further achieve a quintuplet acceleration as well as some feasibility, and efficiency.
@@ 4.8. Set Cover in Graphs
Whereas the register elements are represented by quantum physical subsystems that can store qubits of information, the continuous-time evolution depends explicitly on a Hamiltonian that defines the interactions between register elements (Kaminsky2004, ). An ideal Hamiltonian may allow for arbitrary interactions between elements, but physical and technological limitations often prevent the fabrication of arbitrary interactions or forms of connections in actual devices. A prominent example is found in the quantum annealer developed by D-Wave Systems, Inc. (Johnson2011Quantum, ), which uses a well-defined hardware connectivity graph called the Chimera lattice and implements problems that can be described using the Ising Hamiltonian with two-body interactions Cai, MacCready, and Roy (CMR) have presented a randomized algorithm for generating an embedding (cai2014practical, ). To find the shortest path between randomly mapped logical vertices, their approach is based on Dijkstra’s algorithm. The method is not guaranteed to work and has the worst case complexity, which scales as $\mathcal{O}(n^{9})$ with the input graph order n, although the average case behavior appears to be $\mathcal{O}(n^{3})$. The CMR embedding algorithm represents a significant portion of the time needed for a quantum annealing workflow, and for even modest problem sizes it can far exceed the time required for executing a quantum annealing schedule (Humble_2016, ).
@@ 4.9. Graph Coloring
Consider two separated proves, an interrogator, and a given graph $G(V,E)$. The quantum chromatic number (QN) is the minimal number of colors necessary in a protocol with which the provers may assure the interrogator that they have a coloring of $G$.
@@
Bravyi et al. (bravyi2022hybrid, ) proved a way to apply the Recursive Quantum Approximate Optimization algorithm (RQAOA) to MAX-k-CUT. Namely, the problem of finding an approximate vertex k-coloring of a graph. The authors performed a comparison between the best-known classical and hybrid classical-quantum algorithms. They initially showed how QAOA fails to solve the above optimization problem for most of the regular bipartite graphs at every constant level $p$.The approximation ratio one may obtain by using QAOA hardly outperforms a random assignment of colors to the vertices. As their next step, the authors constructed a classical simulation algorithm to simulate level-1 QAOA and level-1 RQAOA for arbitrary graphs. They had up to $300$ qutrits which were applied to randomly generated 3-colorable constant-degree graphs. The simulation showed that level-1 RQAOA is competitive. That is, for the considered groups of graphs, the approximation ratios were usually higher than the generic classical algorithm based on rounding an SDP relaxation. The authors conjecture that higher-level RQAOA may be a beneficial algorithm for NISQ devices.
@@ 4.10. Bisection Problem
Younes (younes2015bounded, ) introduced a bounded-error quantum polynomial time (BQP) algorithm for the max-bisection and the min-bisection problems. This algorithm runs in $\mathcal{O}(m^{2})$ for a graph with $m$ edges and in the worst case runs in $\mathcal{O}(n^{4})$ for a dense graph with $n$ vertices. The author represents both problems as Boolean constraint satisfaction problems s.t. the set of satisfied constraints is simultaneously maximized/minimized. This is done by an iterative partial negation and partial measurement technique. The algorithm targets a general graph. It also achieved a probability of success $1-\varepsilon$ for $\varepsilon>0$, using polynomial space resources.
@@ 4.11. Finding Cliques in Graphs
Miyamoto et al. (miyamoto2020quantum, ) showed how to apply the combination of quantum search with classical dynamic programming to the minimum Steiner tree problem. The complexity of their quantum algorithm is $\mathcal{O}((1.812)^{k}poly(n))$ s.t. $n$ is the number of vertices in the graph, and $k$ shows the number of terminals (miyamoto2020quantum, ).
@@ 4.12. Graph and String Edit Distance
An important notion of distance for graph-shaped data is the Graph Edit Distance (GED). As the complexity of computing GED is the same as NP-hard problems, it is reasonable to consider approximate solutions. Incudini et al. (Incudini_2022, ) presented a QUBO solution to the GED problem. This gives us two possibilities to implement quantum annealing and variation algorithms running on both types of quantum hardware that exist today: quantum annealer and gate-based quantum computers (Incudini_2022, ). Quantum annealing is a process using quantum fluctuations to optimize for finding the general minimum of an objective function over some set of possible solutions.
The authors in (boroujeni2021approximating, ) introduce a quantum constant approximation algorithm to compute the edit distance in truly subquadratic time. The algorithm’s runtime is proportional to $\mathcal{O}(n^{1.858})$ in terms of quantum computation. It approximates the edit distance by a factor of seven. Also, the authors extended the above result to an $\mathcal{O}(n^{1.781})$ quantum algorithm, approximating the edit distance within a larger constant factor.
@@ 4.13. String-Related Problems
Definition 4.17 (Longest Common Extension(LCE)).
@@
Consider the decisional version of the Longest Common Substring problem (LCS). This variant is provided with a threshold of $d$. It asks whether two length-n input strings have a common substring of length $d$. The two corner cases of $d=1$ and $d=n$, correspond to Element Distinctness and Unstructured Search, respectively.
The authors proved that the complexity of LCS with threshold $d$ smoothly hovers between the two extreme cases up to $n^{o(1)}$ factors. As stated in (jin2023quantum, ), LCS with threshold $d$ has a quantum algorithm with time and query complexity of $n^{2/3+o(1)}/d^{1/6}$. Also, a lower bound of $\Omega(n^{2/3}/d^{1/6})$ holds for the quantum query complexity (jin2023quantum, ). The above results suggest improvement over the previous upper bounds of $\tilde{\mathcal{O}}(\min\{n/d^{1/2},n^{2/3}\})$ obtained in (legall_et_al:LIPIcs.ITCS.2022.97, ; doi:10.1137/1.9781611977073.109, ).
The main contribution in terms of technicality in (jin2023quantum, ) is the quantum speed-up of the Synchronizing Set appeared in (10.1145/3519935.3520061, ). This technique implies a near-optimal LCE data structure under the quantum setting as well.
@@ 4.14. Chemistry and Bio-informatics problems
Consider the problem of computing the number of elements in each hitting-set in an instance of the hitting-set problem. Chang et al. (chang2010quantum, ) showed that the quantum variant of bio-molecular solutions may play the oracle’s role in Grover’s algorithm. Namely, the target state labeling, getting executed before Grover’s searching steps. The authors also performed a three-qubit nuclear magnetic resonance (NMR) experiment to solve the simplest variant of the hitting-set problem.
@@ 4.15. Mixed-Integer Programming
Ajagekar et al. (ajagekar2022hybrid, ) developed QC-based solution strategies, exploiting quantum annealing and classical optimization techniques. This helps solve large-scale scheduling problems in manufacturing systems. They illustrated the applications of their algorithms in two case studies in production scheduling: First, a hybrid QC-based solution for the job-shop scheduling problem is suggested. Second, the authors introduce a hybrid QC-based parametric method for the multipurpose batch scheduling problem with a fractional objective. The algorithms tackle optimization problems formulated in terms of mixed-integer linear and mixed-integer fractional programs, respectively. They also provide feasibility guarantees.
Ajagekar et al. conducted a comparison between the performance of the SOTA solvers (exact and heuristic) and their proposed QC-based hybrid solution. The comparison took place on both job-shop and batch-scheduling problems. The hybrid frameworks suggested by the authors use quantum annealing to support their established deterministic optimization algorithms.
Many approaches to quantum optimization solve unconstrained binary programming problems. Nonetheless, mixed-integer linear programming provides more insight into practice. The authors in (chang2020hybrid, ) developed a new approach for mixed-integer programming. Their approach applies Benders decomposition. This helps decompose the mixed-integer programming into binary programming and linear programming sub-problems. The above sub-problems are solvable by a noisy intermediate-scale quantum processor and a conventional processor, respectively. The algorithm suggested by the authors provably reaches the optimal solution of the original mixed-integer programming problem. It is also tested in practice. In particular, on a D-Wave 2000Q quantum processing unit. The algorithm turned out to be practical for small-scaled test cases. Inspired by power system applications, the authors tested the algorithm on mixed-integer programming as well.

### v4

@@ 5.3. Graph Traversal Problems
One of the most famous problems in graph theory is the travelling salesman problem. Which closely relates to the Hamiltonian Cycle problem. However, the extent to which quantum computers can speed up algorithms to resolve this problem is not yet known. Moylett et al. (moylett2017quantum, ) use a quantum backtracking algorithm to a classical algorithm by Xiao and Nagamochi to demonstrate a quadratic quantum speedup when the degree of each vertex is at least 3. Then, to speed up the classical algorithm for when the degree of each vertex is at least 4, authors use similar techniques to speed up the higher degree graphs by reducing the instances (moylett2017quantum, ).
Salehi et al. (salehi2022unconstrained, ) analyze the travelling salesman problem with time windows (TSPTW) in the context of quantum computation. They introduced quadratic unconstrained and higher-order binary optimization formulations as well as the advantages of edge-based and node-based formulations of TSPTW. In addition, the authors conducted experimental realizations on their methods on a quantum annealing device.
@@ 5.4. Maximum Matching on Graphs
The work in (dorn2009quantum, ) introduces quantum algorithms for the matching problems. The author considers unweighted and weighted graphs with $n$ vertices and $m$ edges. In general graphs, the results in (dorn2009quantum, ) shows that finding a maximal matching takes $\mathcal{O}(\sqrt{nm}\log^{2}n)$. In addition, a maximum matching under the same setting takes $\mathcal{O}(n\sqrt{m}\log^{2}n)$. Considering a different setting, the author proves that finding a maximum weight matching in bipartite graphs is proportional to $\mathcal{O}(n\sqrt{m}N\log^{2}n)$ s.t. $N$ is the largest edge weight. Whilst a minimum weight perfect matching introduces the time complexity of $\mathcal{O}(n\sqrt{nm}\log^{3}n)$.
QA suits problems that can be formulated as QUBO problems. E.g. SAT, graph colouring, and traveling salesman. McLeod and Sasdelli (mcleod2022benchmarking, ) investigate the performance of D-Wave’s 2000Q (2048 qubits) and Advantage (5640 qubits) quantum annealers empirically. In particular, they benchmark the solvers based on solving a specific instance of the maximum cardinality matching problem. They used the results of a previous work that measured the performance of earlier QA hardware from D-Wave. The authors illustrated that Advantage produces optimal solutions to problems of a larger size, compared to 2000Q. Further, the authors investigated the scaling of the diabatic transition probability, by utilizing the Landau-Zener formula. In particular, they considered the structure of the problem and its implications for suitability to QA. Moreover, they introduced a method, exploring the behavior of minimum energy gaps for scalable problems deployed to QAs. The authors showed for their targeted QA problem that the minimum energy gap does not scale as desired. One might ask if this problem is suitable for benchmarking QA hardware. It might lack the necessary nuance to recognize significant performance improvements between generations.
Given an $L\crossproduct L$ array of qubits. Fowler (fowler2013minimum, ) considers the minimum weight perfect matching problem. Provided $(1)$ a modular architecture s.t. there is a finite maximum number of non-functioning qubits in any given connected defective patch $(2)$ gate, leakage, and loss error rates below some set of nonzero values $(3)$ a 2-D array of processors with finite processing power, finite local memory, and a communication line with their nearest neighbour at a finite data rate $(4)$ external memory to store all detection events and matching data. This should last for the duration of a temporally finite quantum algorithm.
@@ 5.5. Maximum-Cut Problem
Boulebnane and Montanaro (boulebnane2021predicting, ) generalized techniques appeared in (farhi2022quantum, ). Applied to different families of sparse random graphs, the authors evaluated the performance of QAOA in the infinite-size limit for the MAX-CUT problem. Their analysis rigorously holds only when $n\rightarrow\infty$ and large degree limit. Still, the analysis provides sufficient initial variational parameters to optimize small-scale, very sparse instances. In addition, the authors introduced a numerical algorithm. It evaluates the performance of QAOA on an average finite-size instance of the SK model. This is where the analytic methods exposed in (farhi2022quantum, ) fail. Nevertheless, the algorithm supports low depth only. Further, their work cannot address the D-spin model for $D\geq 3$ and $p>1$.
@@ 5.6. Graph Isomorphism
Mondada and Andrés-Martínez (mondada2023subgraph, ) suggested an algorithm for the subgraph isomorphism problem on port graphs. Their algorithm performs pattern matching in quantum circuits for many patterns simultaneously. This is independent of the number of patterns. The running time of the algorithm is linear for the size of the input quantum circuit, after pre-computation, in which the patterns are compiled into a decision tree. Consider the connected port graphs. Such graphs introduce a property i.e. every edge $e$ incident to $v$ has a label $L_{v}(e)$ unique in $v$. The algorithm suggested by Mondada and Andrés-Martínez (mondada2023subgraph, ) enumerates the subgraph isomorphism problem $H\subseteq G$ for all $m$ pattern matches in time $\mathcal{O}(P)^{P+\frac{3}{2}}\cdot|V(G)|+\mathcal{O}(m)$ s.t. $P$ is the number of vertices of the largest pattern. Also, they can express the bound obtained with the quantum circuit in terms of the maximum number of qubits $N$ and depth $\delta$ of the patterns: $\mathcal{O}(N)^{N+\frac{1}{2}}\cdot\delta|V(G)|\log(\delta)+\mathcal{O}(m)$.
@@ 5.7. Minimum Vertex Cover Problem
Zhang et al. in (zhang2022applying, ) provided a quantum circuit solution scheme. They used an algorithm of quantum approximate optimization. A quantum Ising equation and Hamiltonians are obtained based on the Ising model quantized by rotation operators and Pauli operators. Also, a parametric unitary transformation is acquired between the original Hamiltonian and the problem Hamiltonian, as a generator. The final quantum state and the problem of Hamilton’s expectation are generated by a continuous evolution of two parametric unitary transformations. The results of simulations in (zhang2022applying, ) indicate that the model may solve problems with large probability in linear time. It can further achieve a quintuplet acceleration as well as some feasibility, and efficiency.
@@ 5.8. Set Cover in Graphs
Whereas the register elements are represented by quantum physical subsystems that can store qubits of information, the continuous-time evolution depends explicitly on a Hamiltonian that defines the interactions between register elements (Kaminsky2004, ). An ideal Hamiltonian may allow for arbitrary interactions between elements, but physical and technological limitations often prevent the fabrication of arbitrary interactions or forms of connections in actual devices. A prominent example is found in the quantum annealer developed by D-Wave Systems, Inc. (Johnson2011Quantum, ), which uses a well-defined hardware connectivity graph called the Chimera lattice and implements problems that can be described using the Ising Hamiltonian with two-body interactions Cai, MacCready, and Roy (CMR) have presented a randomized algorithm for generating an embedding (cai2014practical, ). To find the shortest path between randomly mapped logical vertices, their approach is based on Dijkstra’s algorithm. The method is not guaranteed to work and has the worst case complexity, which scales as $\mathcal{O}(n^{9})$ with the input graph order n, although the average case behaviour appears to be $\mathcal{O}(n^{3})$. The CMR embedding algorithm represents a significant portion of the time needed for a quantum annealing workflow, and for even modest problem sizes it can far exceed the time required for executing a quantum annealing schedule (Humble_2016, ).
@@ 5.9. Graph Coloring
Consider two separated proves, an interrogator, and a given graph $G(V,E)$. The quantum chromatic number (QN) is the minimal number of colours necessary in a protocol with which the provers may assure the interrogator that they have a colouring of $G$.
@@
Bravyi et al. (bravyi2022hybrid, ) proved a way to apply the Recursive Quantum Approximate Optimization algorithm (RQAOA) to MAX-k-CUT. Namely, the problem of finding an approximate vertex k-coloring of a graph. The authors performed a comparison between the best-known classical and hybrid classical-quantum algorithms. They initially showed how QAOA fails to solve the above optimization problem for most of the regular bipartite graphs at every constant level $p$.The approximation ratio one may obtain by using QAOA hardly outperforms a random assignment of colours to the vertices. As their next step, the authors constructed a classical simulation algorithm to simulate level-1 QAOA and level-1 RQAOA for arbitrary graphs. They had up to $300$ qutrits which were applied to randomly generated 3-colorable constant-degree graphs. The simulation showed that level-1 RQAOA is competitive. That is, for the considered groups of graphs, the approximation ratios were usually higher than the generic classical algorithm based on rounding an SDP relaxation. The authors conjecture that higher-level RQAOA may be a beneficial algorithm for NISQ devices.
@@ 5.10. Bisection Problem
Younes (younes2015bounded, ) introduced a bounded-error quantum polynomial time (BQP) algorithm for the max-bisection and the min-bisection problems. This algorithm runs in $\mathcal{O}(m^{2})$ for a graph with $m$ edges and in the worst case runs in $\mathcal{O}(n^{4})$ for a dense graph with $n$ vertices. The author represents both problems as Boolean constraint satisfaction problems s.t. the set of satisfied constraints is simultaneously maximized/minimized. An iterative partial negation and partial measurement technique do this. The algorithm targets a general graph. It also achieved a probability of success $1-\varepsilon$ for $\varepsilon>0$, using polynomial space resources.
@@ 5.11. Finding Cliques in Graphs
Miyamoto et al. (miyamoto2020quantum, ) showed how to apply the combination of quantum search with classical dynamic programming to the minimum Steiner tree problem. The complexity of their quantum algorithm is $\mathcal{O}((1.812)^{k}poly(n))$ s.t. $n$ is the number of vertices in the graph, and $k$ shows the number of terminals.
@@ 5.12. Graph and String Edit Distance
An important notion of distance for graph-shaped data is the Graph Edit Distance (GED). As the complexity of computing GED is the same as NP-hard problems, it is reasonable to consider approximate solutions. Incudini et al. (Incudini_2022, ) presented a QUBO solution to the GED problem. This gives us two possibilities to implement quantum annealing and variation algorithms running on both types of quantum hardware that exist today: quantum annealer and gate-based quantum computers. Quantum annealing is a process using quantum fluctuations to optimize for finding the general minimum of an objective function over some set of possible solutions.
Boroujeni et al. (boroujeni2021approximating, ) introduced a quantum constant approximation algorithm to compute the edit distance in truly subquadratic time. The algorithm’s runtime is proportional to $\mathcal{O}(n^{1.858})$ in terms of quantum computation. It approximates the edit distance by a factor of seven. Also, the authors extended the above result to an $\mathcal{O}(n^{1.781})$ quantum algorithm, approximating the edit distance within a larger constant factor.
@@ 5.13. String-Related Problems
Definition 5.17 (Longest Common Extension(LCE) (a.k.a. Longest Common Subsequence (LCS))).
@@ 5.14. Chemistry and Bio-informatics problems
Consider the problem of computing the number of elements in each hitting-set in an instance of the hitting-set problem. Chang et al. (chang2010quantum, ) showed that the quantum variant of bio-molecular solutions may play the oracle’s role in Grover’s algorithm. Namely, the target state labelling, getting executed before Grover’s searching steps. The authors also performed a three-qubit nuclear magnetic resonance (NMR) experiment to solve the simplest variant of the hitting-set problem.
@@ 5.15. Mixed-Integer Programming
Ajagekar et al. (ajagekar2022hybrid, ) developed QC-based solution strategies, exploiting quantum annealing and classical optimization techniques. This helps solve large-scale scheduling problems in manufacturing systems. They illustrated the applications of their algorithms in two case studies in production scheduling: First, a hybrid QC-based solution for the job-shop scheduling problem is suggested. Second, the authors introduce a hybrid QC-based parametric method for the multipurpose batch scheduling problem with a fractional objective. The algorithms tackle optimization problems formulated in terms of mixed-integer linear and mixed-integer fractional programs, respectively. They also provide feasibility guarantees. Ajagekar et al. conducted a comparison between the performance of the SOTA solvers (exact and heuristic) and their proposed QC-based hybrid solution. The comparison took place on both job-shop and batch-scheduling problems. The hybrid frameworks suggested by the authors use quantum annealing to support their established deterministic optimization algorithms.
Many approaches to quantum optimization solve unconstrained binary programming problems. Nonetheless, mixed-integer linear programming provides more insight into practice. Chang et al. (chang2020hybrid, ) developed a new approach for mixed-integer programming. Their approach applies Benders decomposition. This helps decompose the mixed-integer programming into binary programming and linear programming sub-problems. The above sub-problems are solvable by a noisy intermediate-scale quantum processor and a conventional processor, respectively. The algorithm suggested by the authors provably reaches the optimal solution of the original mixed-integer programming problem. It is also tested in practice. In particular, on a D-Wave 2000Q quantum processing unit. The algorithm turned out to be practical for small-scaled test cases. Inspired by power system applications, the authors tested the algorithm on mixed-integer programming as well.